# Unisolver: PDE-Conditional Transformers Towards Universal Neural PDE Solvers

### v4

@@ I.7 Detailed Compute Resources
This paper presents Unisolver to solve PDEs under universal PDE components, which achieves impressive performance supported by extensive analyses and visualizations. However, our method is currently limited to grid data due to the patchifying process during the embedding of point-wise components. Actually, this limitation is shared in all the generalizable PDE solvers, such as MPP [29], Poseidon [13], PDEformer [45] and DPOT [11]. One fundamental reason is the lack of suitable and large-scale irregular-mesh PDE datasets, which will require extremely high computation costs for generation and massive resources for collection. Since our primary focus in this paper is on the study of model architecture design and generalization capabilities, we would like to leave the irregular-mesh PDE dataset as a future work. Also, the capability to handle irregular meshes of Unisolver can be achieved by replacing the canonical Transformer with the latest geometry-general PDE solver: Transolver [43].

### v5

@@ I.7 Detailed Compute Resources
This paper presents Unisolver to solve PDEs under universal PDE components, which achieves impressive performance supported by extensive analyses and visualizations. However, our method is currently limited to grid data due to the patchifying process during the embedding of point-wise components. Actually, this limitation is shared in all the generalizable PDE solvers, such as MPP (McCabe et al., 2023), Poseidon (Herde et al., 2024), PDEformer (Ye et al., 2024) and DPOT (Hao et al., 2024). One fundamental reason is the lack of suitable and large-scale irregular-mesh PDE datasets, which will require extremely high computation costs for generation and massive resources for collection. Since our primary focus in this paper is on the study of model architecture design and generalization capabilities, we would like to leave the irregular-mesh PDE dataset as a future work. Also, the capability to handle irregular meshes of Unisolver can be achieved by replacing the canonical Transformer with the latest geometry-general PDE solver: Transolver (Wu et al., 2024).